# Dynamic read & write optimizaton with TurtleKV

### v1

@@ 1. Introduction
Key-value storage is a foundational building block upon which many data-management applications are built. Though key-value storage’s dictionary-based interface flexibly supports a diverse range of applications, concrete implementations must trade off between the cost of reads, writes, and memory usage. Prior work hypothesizes that only two of these dimensions can be optimized at the cost of the third (Brodal and Fagerberg, 2003; Athanassoulis et al., 2016). The result is a proliferation of implementations tuned for specific workload characteristics (Meta Platforms, Inc., 2022; Conway et al., 2020; Conway et al., 2023; Apache, 2025; Google, [n.d.]). This growing diversity complicates database management: organizations must often deploy multiple databases to support different workloads and rely on expensive Extract-Transform-Load (ETL) pipelines to move data (Databricks, [n.d.]).
Figure 1. Today’s key-value engines (left) trade reads directly against writes, then compensate for lost performance by spending memory. In contrast, TurtleKV(right) uses a balanced on-disk data structure. It dynamically trades off memory for increased write or read performance.
With respect to the tradeoffs, state-of-the art databases limit memory usage and choose between reads or writes for the remaining dimension they wish to optimize (left diagram in Figure 1). They use on-disk structures biased for performance in this remaining dimension and use other mechanisms to recover lost performance in the unoptimized dimension. For example, SplinterDB (Conway et al., 2020; Conway et al., 2023) uses a on-disk data structure, STBϵ-trees, biased toward writes and uses a limited amount of memory for quotient filters/maplets to improve read performance. TreeLine (Yu et al., 2022) uses update-in-place grouping mechanisms and spends memory on histograms to improve reads. Industry approaches, such as RocksDB (Meta Platforms, Inc., 2022), optimize for writes via LSM-trees and use Bloom filters to improve read performance. These approaches are static. They find a single read/write optimization point and permanently change the on-disk data structure, creating a residual memory cost for the entire lifetime of the data.
Though powerful, the above approaches curtail the maximum achievable performance in the unoptimized dimension. As a consequence, supporting load-then-read workloads with high performance requires either spending significant I/O changing the on-disk format or using an expensive ETL pipeline to move data between different databases.
We present TurtleKV, a key-value engine that uses an unbiased on-disk structure suited for reads and writes. It includes a tuning knob that dynamically controls the allocation of memory resources on demand within a fixed budget to improve read or write performance as needed (right diagram in Figure 1). The knob allows memory usage to be re-allocated to the alternate dimension (e.g., reads) when high performance in currently-chosen dimension is no longer needed (e.g., writes). TurtleKV is made possible by the following two observations. First, though the trade-off between memory, reads and writes may be fundamental, the choice of which dimensions to trade-off against are not. Second, many common workloads do not need high write and read performance simultaneously, so memory can be fungibly exchanged between these two dimensions as needed. For workloads that require high read-and-write performance simultaneously, TurtleKV allows human operators to adjust the tuning knob in a weighted way to reflect their operational goals.
To support TurtleKV, we introduce a new external memory dictionary data structure, the TurtleTree, which improves read performance and bounds worst case write I/O behavior. TurtleTrees are a fusion of B+-trees and LSM-trees, leveraging the strengths of both to achieve an unbiased state, balanced between read and write performance. B+-trees, which offer high read performance, provide the high-level structure with data organized in tree-node pages. LSM-trees, which offer higher write performance, are used within the tree-nodes to buffer updates. By carefully managing the TurtleTree structure, we can control the rate at which updates flow down the tree, bounding the number of I/Os required for reads and writes. This property results in a balanced on-disk structure that can be dynamically tuned for greater read or write performance.
We identify TurtleTrees as belonging to a previously unidentified category of data structures, which we name Bϵ+-trees. Such Bϵ+-trees differ from Bϵ-trees and other buffered B-trees by adding a level of indirection between tree-node pages and pages containing per-node buffered updates. SplinterDB’s size-tiered Bϵ-trees (STBϵ-trees) also belong to the Bϵ+-tree family. TurtleTrees differ from STBϵ-trees in that they use level-tiered LSM-trees instead of size-tiered ones. This change shifts the focus from write performance to more unbiased point between reads and writes, while also allowing for more fine-grained control over write amplification. In contrast to traditional level-tiered LSM-trees, TurtleTree’s design limits write I/O costs by bounding the number of pages flushed to disk at any time. Furthermore, TurtleTrees and other data structures in the Bϵ+-tree family are complementary to existing read/memory trade-off mechanisms such as page caching and AMQ filtering because of the separation of data (leaf pages and buffer segments) from metadata (interior nodes and any filters).
Checkpoint distance ($\chi$), which is TurtleKV’s dynamic tuning knob, controls the portion of TurtleKV’s total memory devoted to improving write performance and can achieve the performance of write-optimized, size-tiered approaches on demand. Concretely, this knob sets the number of in-memory batch updates that can be performed on a TurtleTree before updates are written to disk and combines synergistically with TurtleTree ’s control over flushed pages. A higher value allows more data to be merged in memory, reducing write amplification and thus increasing write performance. To dynamically improve read performance when under cache pressure, TurtleKV can choose a smaller value, so that memory that is unused for writes maybe (re-)allocated to cache more data pages as needed. The key benefit of this knob is that unlike existing systems that permanently spend memory by creating new structures such as filters, tuning decisions are not anchored to on-disk structures at the point data is written. TurtleKV can dynamically optimize reads or writes by simply adjusting the allocation of memory.

### v2

@@ 1. Introduction
Today the builders of complex, cloud-based applications and systems have many different key-value stores to choose from, each with their own biases and trade-offs (e.g., read-optimized vs write-optimized; online-vs-analytic, in- vs out-of-memory, ordered-vs-hashed, domain-specific-vs-general-purpose, etc.)(jermaineNovelIndexSupporting1999; chaBlinkHashAdaptive2023; giladEvenDBOptimizingKeyvalue2020; pelkonenGorillaFastScalable2015; ashkianiGPULSMDynamic2018; wangLavaStoreByteDancesPurposeBuilt2024; sollezaMachPluggableMetrics2022; renSlimDBSpaceefficientKeyvalue2017a). These are frequently difficult to tune to accommodate evolving requirements after a system is built (RocksDBTuningGuide; caoCharacterizingModelingBenchmarking; SmallDatumImpact2022; jagadeesanWorstCaseAverageCaseAnalysis2020; RevisitingB+treeVs2022; nguyenCharacteristicsMongoDBTrimbased2016; andreoliInducingHugeTail2023; fedorovaPerformanceComprehensionWiredTiger2018). As a consequence, we see a proliferation of multi-stores with complex and expensive logic for synchronization, data transfer, and invariant maintenance (etl_pipeline; kumarValuePropositionETL; mukherjeeComparativeReviewData2017; el-sappaghProposedModelData2011; limSILTMemoryefficientHighperformance). Engineering effort must be spent multiple times for development, debugging, optimization, and maintenance of each storage engine.
One would not be without reason believing that a multitude of storage engines is inevitable. The RUM Conjecture (athanassoulisDesigningAccessMethods2016) posits a fundamental, three-way trade-off in the design of data storage/access methods between the resource overhead of reads (R), updates/writes (U), and memory/storage (M). If correct, this might seem to imply that the divergence of requirements for different workloads and applications means key-value stores inhabit a problem space where one size does not, and cannot, fit all(idreosKeyValueStorageEngines2020). Our work assumes the correctness of the RUM Conjecture as its starting point, but seeks to challenge this implication and propose an alternative approach to key-value storage engine design in the face of inherent trade-offs.
The notion of a three-way trade-off in the RUM Conjecture means that at least two trade-off selection mechanisms, or tuning knobs, must be utilized to fully cover the space of optimal possibilities; e.g. reads-vs-writes (RW), reads-vs-memory (RM), or writes-vs-memory (WM). Our key insight is that these basic trade-offs have important differences, and that these differences make some inherently better suited for exploitation as tuning knobs. For example, if a suboptimal tuning point is selected for the read-memory trade-off, it is likely easy to correct because of the mechanisms involved: simply modify the cache size to decrease either the I/O cost of queries or the memory footprint of the system at the expense of the other. In contrast, if a system is improperly tuned with respect to the read-write trade-off, things aren’t so simple. This is because read-write tuning is largely accomplished via selection of primary data structure, which affects implementation (code) and the invariants of stored data, both of which are expensive to modify post-facto. While some dictionary data structures do support parameter-based read-write trade-off tuning (e.g., compaction policy in LSM-trees with adjustable fanout (oneilLogstructuredMergetreeLSMtree1996; luoLSMbasedStorageTechniques2020) ), when analyzed closely and evaluated empirically, we find the mechanisms in common use to be sub-optimal. In any case, even if the primary data structure of a system does support smooth, optimal tuning between read and write optimization (e.g, Bϵ-trees’ $\epsilon$ parameter (brodalLowerBoundsExternal2003; benderIntroductionBetreesWriteOptimization2015)), bias is still encoded at the storage level, making re-tuning expensive.
Since write-optimization is very important to certain workloads (caoCharacterizingModelingBenchmarking), write-memory (WM) tuning must be supported as a first-class feature if we wish to avoid the downsides of RW tuning bias in stored data structures. To do this it is clearly necessary, though insufficient, to simply expose a tuning knob. The choice of data structure and the overall design of the system must work together to maximize the effectiveness of this trade-off. A good example of how data structures can anticipate a resource trade-off (RM) in a system design context is the pervasive use of B+-trees in preference to B-trees (comerUbiquitousBTree1979; bayerOrganizationMaintenanceLarge1970). By storing all key-value records at the bottom level of the tree, B+-trees offer increased branching factor and the possibility for all interior nodes to be cached in memory, reducing read amplification to $O(1)$. A key idea of the work presented here is to apply a similar principle in data structure selection to enhance the effectiveness of the write-memory (WM) trade-off, elevating it to first-class status in key-value system design.
We present TurtleKV, a proof-of-concept for a new approach to key-value storage oriented around RM and WM tuning knobs which dynamically navigate the RUM trade-off space without having to change the structure of stored data. TurtleKV is built around a novel data structure, the TurtleTree, which offers two important properties: 1. it achieves a neutral, balanced compromise between the extremes of read- and write-optimized external memory dictionaries, and 2. it affords a smooth trade-off between memory usage and write amplification which is simultaneously efficient, scale-independent (in terms of data size $N$), conceptually simple, and wide-ranged in effect. TurtleKV implements a system architecture which is similar at a high level to conventional LSM-based key-value stores, with the addition of several crucial optimizations to eliminate in-memory bottlenecks so that reductions in write-amplification translate to increased throughput. Because TurtleKV starts from a fundamentally stronger premise, it avoids many of the complex workarounds and features which tend to creep into modern key-value stores in order to try to mitigate their intrinsic limitations. As a result, we show that TurtleKV outperforms today’s key-value engines, while offering superior tunability, with a modest code base of similar or lesser complexity.
In this work we demonstrate that write-vs-memory tuning mechanisms are often not given first-class status in the design of current key-value stores ($\S$2.2). For example, while RocksDB (RocksDBPersistentKeyvalue), one of the most popular key-value storage engines today, supports runtime adjustment of write buffer size (RocksDBTuningGuide), and our experiments show this parameter is highly effective at spending memory to reduce write amplification, doing so yields very little benefit in terms of increased end-to-end performance. WiredTiger (WiredTigerWiredTigerDeveloper), another popular storage engine on the opposite end of the read-write spectrum, also supports WM tuning, and even demonstrates end-to-end benefit in terms of increased insert throughput over wide range of tuning points. However, WiredTiger’s B+-tree based design offers little benefit at low-to-moderate memory allocations due to the inherent limitations of the data structure (idreosDesignContinuumsPath2019). Finally, we find that SplinterDB (conwaySplinterDBClosingBandwidth2020; conwaySplinterDBMapletsImproving2023; SplinterDBb) which is optimized for high concurrency and fast storage devices, is able to take advantage of higher memory allocations to reduce write amplification and average insertion latency for randomized insertion workloads. However, since SplinterDB does not offer any way to express a preference for optimization target (i.e., use cache memory to optimize writes vs reads), in mixed workloads the optimization effect of higher cache size may become unpredictable, as the dualing concerns of updates and queries vie for limited memory resources.
The contributions of this work are:
  1. (1)
    An analysis of how current key-value stores and their data structures fall short of fully exploiting the potential of the RUM space, especially with respect to the write-memory trade-off
  2. (2)
    A novel external memory dictionary type, the TurtleTree, within a previously unidentified class of data structures, Bϵ+-trees, which combines ideas from both read- and write-optimized dictionary types
  3. (3)
    A key-value storage engine design, TurtleKV, which uses TurtleTrees to offer simple and smooth trade-off tuning for a wide range of resource allocations
  4. (4)
    An experimental evaluation of TurtleKV as compared to SOTA key-value stores WiredTiger, RocksDB, and SplinterDB
2. Read, Update, Memory: Pick Two
Figure 1. RUM as a three-dimensional resource space
The RUM Conjecture (athanassoulisDesigningAccessMethods2016) states that if one holds any of the three resource overheads (read, update/write, and memory/space) constant, the remaining two are placed in a zero-sum, trade-off relationship. Geometrically, this can be visualized, as in Figure 1, by a three dimensional resource space, where points represent the actual resource cost of a concrete data access method with respect to a particular workload. The notion of a three-way trade-off, in this conceptualization, is represented by a two dimensional surface or manifold embedded within the larger-dimensional space (dark gray). Fixed resource commitments can be thought of as two dimensional slices of this surface orthogonal to the axis of the fixed resource, with the resulting intersection defining the trade-off curve between the remaining two resources.
Navigating the optional two-dimensional subset of the RUM space requires the use of exactly two mechanisms, or tuning knobs, which serve as a coordinate system within the sub-space. In this section we analyze our problem space by considering three tuning knob candidates: the read-write trade-off (RW), the read-memory trade-off (RM), and the write-memory trade-off (WM). These are explored through the perspectives of data structure ($\S$2.1) and system design ($\S$2.2). We conclude with a sketch of the key ideas behind TurtleKV ($\S$2.3).
2.1. Data Structures: Trade-off Tuning
The most consequential design decision for a key-value storage engine is its choice of primary data structure. Here we review the major categories of external memory dictionary, grouped by their RW trade-off optimization target (read-optimized, write-optimized, and hybrid/balanced). We highlight the essential insights of each data structure, how each approaches trade-off tuning, and their inherent limitations.
2.1.1. Read-Optimized (B-trees).
B-trees(comerUbiquitousBTree1979) provide optimal (yiDynamicIndexabilityOptimality2012) read performance and can utilize memory to reduce the number of reads required to process queries by caching pages containing tree-nodes (RM trade-off). B+-trees, the most commonly implemented B-tree variant, exploit this mechanism by placing key/value data in the leaves, allowing all but the bottom layer of the tree to be cached in main memory so that queries can be answered with a single read.
State-of-the-art approaches to reducing the write cost of updates in B-trees employ the use of a write-ahead log (WAL) to ensure durability of updates, plus a scheme for applying updates to the tree in memory (WM trade-off). We refer to these as Log Structured B-trees. Their goal is to defer rewriting pages containing updated keys, in hopes of collecting enough updates per page to amortize I/O. Bw-trees (bwtrees) do this through the use of delta records applied in-memory to the main tree using lock-/latch-free techniques. The closely related Bf-tree (bftree) uses an alternate strategy of caching and updating sub-page slices (mini-pages) instead full pages, which has additional benefits in terms of lowering the memory cost of caching ”hot” records.
Bw-trees and Bf-trees achieve good write efficiency either when memory is plentiful or when updates are clustered in a few sub-ranges of the key space. In the general case, however, Log Structured B-trees force a sub-optimal trade-off between memory and write amplification, especially as the total data size $N$ grows with respect to available memory $M$ (idreosDesignContinuumsPath2019).
2.1.2. Write-Optimized (LSM-trees)
Log-Structured Merge (LSM) trees (oneilLogstructuredMergetreeLSMtree1996) are a popular write-optimized alternative to B-trees. By deferring the merge (or compaction) of new updates with existing data, they maximize the amount of effective work achieved by each write operation. This results in a dramatic reduction in write amplification, but presents two challenges: first, it increases read costs (and storage footprint) by introducing the possibility that updates for a given key may be stored in multiple locations. Second, since the increasingly high bandwidth of modern storage devices favors performing merges in large sequential runs to maximize I/O utilization, update costs in LSM-trees tend to be amortized costs; i.e. tail latencies can be quite high. This motivates the need for effective de-amortized, or incremental, approaches to compaction.
Page caching is less effective as an RM tuning mechanism for LSM-trees than B-trees because there is potentially much more data to cache. In addition to caching, two primary tuning mechanisms are employed by LSM-tree systems: compaction policy and approximate membership query (AMQ) filtering. Compaction policies divide into greedy (level-tiering) and lazy (size-tiering). In a level-tiered LSM-tree, write cost is increased by repeatedly merging data at the same level (up to $F$ times) to reduce future read cost. In a size-tiered LSM-tree, a tunable number $T$ of key duplicates are allowed to co-exist at a given level before their eventual merge into a larger level. Compaction policy thus alters the invariants of stored data to achieve a RW trade-off mechanism.
Approximate member query (AMQ) filters, e.g. Bloom Filters (bloomSpaceTimeTradeoffs1970), are an RM trade-off mechanism that reduce the cost of gets by having a dedicated filter (or part of a shared filter (conwaySplinterDBMapletsImproving2023)) devoted to each location where a key might reside. AMQ filters offer a tuning parameter $\alpha$ [^1] Also frequently named as $\epsilon$, the false-positive rate, which trades size against accuracy. [^2] A filter must always answer accurately if a query key is not in its set, but it may give a false positive result at rate $\alpha$. The smaller $\alpha$ gets, the more bits per key are required to store the filter. To answer a point query for key $k$, the filters for all locations $S$ where $k$ might reside are consulted before performing any I/O. This rules out, with probability $\prod_{i\in S}{\left(1-\alpha_{i}\right)}$, locations where the query will fail, preventing excess reads.
If filters are used to maintain a target read cost in a size-tiered LSM-tree (with $T$ tiers per level), the total size of the filters must scale in proportion to $T$, since the total number of false-positives for a query is the sum of false-positives for all filters that pertain to a given key (dayanOptimalBloomFilters2018a). As $T$ increases, read amplification increases while write amplification goes down; as filter bits-per-key increases, memory usage rises while read amplification falls due to lower false-positive rate $\alpha$. It is theoretically possible therefore to coordinate the motion of these two parameters ($T,\alpha$) to target a fixed level of read amplification, while writes are effectively traded against memory (WM) using filter size-vs-accuracy as an intermediary. We refer to this strategy as filter arbitrage.
2.1.3. Hybrid-Optimized (Bϵ-trees)
Bϵ-trees (benderIntroductionBetreesWriteOptimization2015; brodalLowerBoundsExternal2003) are a tunable external memory dictionary, able to trade reads against writes. The basic B-tree structure is used as the foundation for a Bϵ-tree. The $\epsilon$ parameter varies the portion of a node page used for pivots, which optimize for reading, and update buffering, which optimizes for writing. At either extreme of their tuning range $\epsilon\in{0..1}$, Bϵ-trees simply degrade into the equivalent of a B-tree or LSM-tree. At $\epsilon=\frac{1}{2}$, the branching factor is $\sqrt{B}$, making tree height twice that of an equivalent B-tree. In this configuration, updates can be flushed down one level at a time from the root in batches of size $\Theta(\sqrt{B})$, dramatically lowering the worst-case write cost and providing a good balanced tuning point.
Figure 2. Bϵ-tree $\rightarrow$ Bϵ+-tree structural transformation
Size-Tiered Bϵ-trees (STBϵ-trees), used by SplinterDB (conwaySplinterDBClosingBandwidth2020; conwaySplinterDBMapletsImproving2023), are a recent re-imagining of the Bϵ-tree designed to close the ”bandwidth gap” between the capabilities of fast NVMe SSDs and traditional key/value storage engine designs. The main structural modification in STBϵ-trees is to add level of indirection between the tree of interior node pages, or trunk, and buffered data, or branches (Figure 2, 4 ). This separation allows buffered updates to be flushed in larger batches than a traditional Bϵ-tree, enabling high I/O bandwidth utilization. It also enables flushes to be decoupled from compactions since only the branch references need to be changed, enabling SplinterDB’s novel flush-then-compact policy. Flush-then-compact has the added benefit of allowing updates with skewed key distributions to be compacted at a lower level of the tree, reducing write amplification for ”hot” records. However, due to their lazy compaction policy, STBϵ-trees must do more work to perform scan queries, which hurts performance.
2.2. System Design: Write-Memory Tuning
We now shift our attention from data structures to the concerns of system design, to explore how well current key-value storage engines support write-vs-memory (WM) trade-off tuning. We present three case studies of systems spanning the read-write (RW) spectrum covered in $\S$2.1: WiredTiger, RocksDB, and SplinterDB. All are high-performance key-value storage engines optimized for concurrent workloads run on hardware with multiple processor cores and solid-state storage devices. RocksDB and WiredTiger in particular are heavily relied upon in real-world systems, and have been the beneficiaries of a large investment in engineering effort. The main differentiator in these systems is the choice of primary data structure: WiredTiger uses B+-Trees [^3] WiredTiger also supports LSM-based indexing, but it is not the default., RocksDB is based on LSM-Trees, and SplinterDB introduces the novel STBϵ-tree.
An investigation into the limitations of how each of these systems support WM trade-off tuning elucidates some of the central issues and provides motivation for the design of our prototype, TurtleKV. Specifically, we present three findings: first, a system’s ability to navigate the WM trade-off is highly dependent on the choice of primary data structure; second, the ability to reduce write amplification does not always translate into increased real-world performance; and third, supporting I/O vs memory trade-offs in general does not imply the ability to target specific tuning points in terms of memory-based read- or write-optimization.
We summarize the key results of these case studies in Table 1.
WiredTiger RocksDB SplinterDB TurtleKV
Data Structure supports
efficient WM trade-off?
System Design has no
in-memory bottlenecks?
First-class write-memory
trade-off tuning knob?
Table 1. The three key requirements for dynamic write-memory trade-off tuning in key-value storage engine design.
2.2.1. WiredTiger
WiredTiger is the default storage engine for MongoDB. It is designed to support high levels of concurrency for transactional key-value workloads. To accomplish this, it uses lock/latch-free algorithms and multi-version concurrency control (avoiding update-in-place) to implement its primary data structure, a B+-tree.
Updates to the B+-tree in WiredTiger take place in ”dirty” buffers pinned to its in-memory page cache, where new versions of existing pages (or entirely new pages) are built. Dirty pages are flushed to durable storage in response to a checkpoint request or when their total size exceeds a runtime-configurable threshold444eviction_dirty_target for background threads, eviction_dirty_trigger for application/foreground threads.. Write-ahead logging can be enabled for durability of record-level updates in between dirty page flush triggers. Thus, WiredTiger supports dynamic write/memory trade-off tuning through its dirty page eviction (write-back) caching mechanism.
Figure 3a shows how WiredTiger responds to increased memory allocation for dirty page buffers, running a workload with uniformly distributed keys inserted in random order. Here we see that average insert latency tracks with write amplification through the entire dynamic range of the resource trade-off, which is good. Reductions in write amplification factor (WAF), however, are minimal for 16mb $\leq M\leq$ 500mb; at higher memory allocations, WAF reduction accelerates. This is because, for random (non-skewed) key distributions, the expected benefit of buffering in B+-Trees is related to the total dictionary size $N$. If each page can hold $B$ items, then the expected write amplification $W_{x}$ for record $x$ is $O(\max(1,\min(\frac{N}{M},B)))$. Note, $W_{x}\in\{1,B\}$, since each record must be written at least once ($1$), and in the worst case it is re-written every time each other record in the same page is written ($B$).
2.2.2. RocksDB
RocksDB is used at Meta as the primary storage engine of ZippyDB (a distributed key/value storage service developed by and used within Meta), UDB (a MySQL storage layer for social graph data) (caoCharacterizingModelingBenchmarking), and by countless other open source projects and commercial systems.
RockDB’s Tuning Guide (RocksDBTuningGuide) presents a formidable set of user-selectable policies and tuning knobs which optimize for various use cases. We identify two such knobs which appear to have the greatest impact on the WM trade-off: the size $\overline{M_{w}}$ of an in-memory table (MemTable) 555rocksdb.options.write_buffer_size, and the number of MemTables $|M_{w}|$666rocksdb.options.min_write_buffer_number_to_merge, rocksdb.options.max_write_buffer_number collected in memory before being flushed to the first level of external storage ($L_{0}$). We find that varying $\overline{M_{w}}$ with fixed $|M_{w}|$ is superior to varying $|M_{w}|$ for fixed size $\overline{M_{w}}$, due to the degradation of query performance as the number of MemTables increases; both approaches offer the same benefit in write-optimization.
RocksDB’s default compaction strategy (leveled) uses incremental merging, in which levels are segmented into fixed-size sorted string table (SSTable) files. When the total size of all segments within a level $L_{i}$ grows too large, an SSTable (of size $\overline{M_{w}}$) within $L_{i}$ is selected to be compacted with data in the next larger/older level $L_{i+1}$. The compaction merges the single segment from $L_{i}$ with all segments in $L_{i+1}$ whose key range overlaps, producing a new sequence of SSTable segments whose keys are ordered and unique. The total write amplification $W_{x}$ for a given record $x$ can be calculated as the number of times $F$ (for fanout) the record is merged into a segment at a given level, times the number of levels in the LSM [^7] The reason there are $\log_{F}({N/M})=\log_{F}{N}-\log_{F}{M}$ levels for data set size $N$ is that the smallest level’s size is always at least $M$; one can think of the first $\log_{F}{M}$ merges of a record as happening in main memory, inside a MemTable.: $W_{x}=F\cdot\log_{F}({N/\overline{M_{w}}})$. Increasing MemTable size $M$ therefore spends memory to reduce write amplification, independent of total size $N$: $\Delta W_{x}=-O(\log{\overline{M_{w}}})$.
(a) WiredTiger
(b) RocksDB
(c) TurtleKV